# Underload: Defending against Latency Attacks for Object Detectors on Edge Devices

### v1

@@ Abstract
Object detection is a fundamental enabler for many real-time downstream applications such as autonomous driving, augmented reality and supply chain management. However, the algorithmic backbone of neural networks is brittle to imperceptible perturbations in the system inputs, which were generally known as misclassifying attacks. By targeting the real-time processing capability, a new class of latency attacks are reported recently. They exploit new attack surfaces in object detectors by creating a computational bottleneck in the post-processing module, that leads to cascading failure and puts the real-time downstream tasks at risks. In this work, we take an initial attempt to defend against this attack via background-attentive adversarial training that is also cognizant of the underlying hardware capabilities. We first draw system-level connections between latency attack and hardware capacity across heterogeneous GPU devices. Based on the particular adversarial behaviors, we utilize objectness loss as a proxy and build background attention into the adversarial training pipeline, and achieve a reasonable balance between clean and robust accuracy. The extensive experiments demonstrate the defense effectiveness of restoring real-time processing capability from $13$ FPS to $43$ FPS on Jetson Orin NX, with a better trade-off between the clean and robust accuracy.
@@ 1 Introduction
Real-time object detection lies at the heart of a wide range of downstream applications such as autonomous driving, drone navigation, video surveillance and augmented reality. From the early detectors such as (Faster) RCNN [1, 2] to the latest versions of the YOLO family [3, 4, 5], we have seen incredible achievements of performance-efficiency improvements, e.g., reaching 56-60% mAP on the MS-COCO benchmark with more than 30 FPS on embedded NVIDIA Jetson boards.
@@ 3.1 Algorithmic Vulnerability from NMS
$T_{\text{nms}}=\left\{\begin{aligned} T_{\text{base}},&~{}~{}~{}|C|\leq N_{t}\\ a|C|^{2},&~{}~{}~{}|C|>N_{t}\\ \end{aligned}\right.$ (1)
@@ 3.2 System-level Impacts of Latency Attacks
$T_{\text{nms}}=(\alpha\frac{|C|^{2}}{S_{\text{IoU}}}+\beta\frac{|C|}{B})<T-T_{\text{backbone}}\$ (3)
$\mathbb{E}[f_{\theta}(x)]_{\text{box}}<\frac{S_{\text{IoU}}}{2\alpha}\sqrt{\frac{\beta^{2}}{B^{2}}-4\frac{\alpha}{S_{\text{IoU}}}(T-T_{\text{backbone}})}$ (4)
@@ 4 Background-Attentive Adversarial Training
$\theta^{\ast}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{(x,y)\sim\mathbb{D}}\bigl{[}\max_{\left\lVert\delta\right\rVert_{p}\leq\epsilon}\mathcal{L}(f_{\theta}(x+\delta),y)\bigr{]}.$ (5)
@@ 4.1 Objectness Loss
$\theta$ $=$ $\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}\sim\mathcal{D},\mathbf{y},\mathbf{b}}\mathcal{L}(f_{\theta}(\mathbf{x}),\mathbf{y},\mathbf{b})$ (6)
@@
$\vspace{-0.03in}\mathcal{L}_{obj}=\sum_{i=1}^{K}\big{[}\mathds{1}_{i}\ell_{BCE}(1,\hat{C}_{i})+(1-\mathds{1}_{i})\ell_{BCE}(0,\hat{C}_{i})\big{]}.\vspace{-0.03in}$ (7)
@@
An essential function of objectness is to differentiate objects from the background. Denote the objects as $x_{\text{obj}}=\bigcup_{k=1}^{K}(b_{i})^{\mathbf{y}}$, $x^{\prime}_{\text{obj}}=x_{\text{obj}}+\delta_{\text{obj}}$, and the background as $x_{\text{bg}}=x-x_{\text{obj}},x^{\prime}_{\text{bg}}=x_{\text{bg}}+\delta_{\text{bg}}$. We define background boundary margin as the distance for a background region to the decision boundary of becoming a “phantom object” in the pixel space $\mathcal{X}$,
@@
$\mathbb{E}_{x\sim\mathbb{D}}[B_{\theta}(x^{\prime}_{\text{bg}})]\leq\mathbb{E}_{x\sim\mathbb{D}}[B_{\theta}(x^{\prime}_{\text{obj}})],$ (9)
@@ 4.2 Proposed AT Method
$\theta^{\ast}=\operatorname*{arg\,min}\limits_{\theta,{\delta^{M}\in\mathcal{S}}}\mathcal{L}\big{(}f_{\theta}(\mathbf{x}+\delta^{M}),\mathbf{b},\mathbf{y}\big{)}\vspace{-0.05in}$ (10)
@@
$\mathcal{S}=\bigl{\{}\delta^{M}\leftarrow~{}\mathbf{M}\odot\operatorname*{arg\,max}\limits_{\left\lVert\delta\right\rVert_{p}\leq\epsilon,i\in\mathcal{K}}\mathcal{L}_{obj}(f_{\theta}(\mathbf{x}+\delta,b_{i},y_{i})\bigr{\}}$ (11)
$\mathbb{E}[f_{\theta}(x+\delta^{M})]_{\text{box}}<\frac{S_{\text{IoU}}}{2\alpha}\sqrt{\frac{\beta^{2}}{B^{2}}-4\frac{\alpha}{S_{\text{IoU}}}(T-T_{\text{backbone}})}\$ (12)
$\mathbf{M}=\{\mathbf{0},\mathbf{1}\}^{r_{x}^{i}\times r_{y}^{i}},r_{x}^{i}\in[0,W_{b}^{i}],r_{y}^{i}\in[0,H_{b}^{i}],i\in\mathcal{K}.$ (13)
@@
Property 3 (Monotoncity). The box count $\mathbb{E}[f_{\theta}(x+\delta^{M})]_{\text{box}}$ is monotonously increasing regarding the mask size $r_{x}^{i}\times r_{y}^{i}$.
@@ 5.5 Background and Object Semantics
Finally, we expose intriguing artifacts between the background and object semantics through the lens of targeted attacks, that echo with our design of background-attentive AT. The goal is to launch targeted attacks to generate phantom objects of “cars” and “boats” under different background regions to different images in Fig. 8.
“#1: a racing car on the track”, “#2: a person on the beach”, “#3: an aeroplane in the sky”, and “#4: a black image with no information”. For #1, it is much easier to generate more “cars” on the track than “boats” and the opposite is true for #2. For #3, it is a bit easier to generate more “boats” in the sky than “cars”, due to the blue color of the sky similar to the ocean. For #4, it is equally possible to generate as many “cars” or “boats” since the blank image contains no semantic information.
From the attacker’s perspective, the attack success rate can be improved by targeting the closest semantics to the background region, e.g., generating phantom “birds” in the “sky”, or “persons” on the “road”, which not only requires less perturbation (stealthier), but also generates more phantoms with higher latency in the NMS. This is because the semantically close objects/background are more vulnerable to adversarial attacks that slightly push the objects/background region over the manifold to be adversarial. From the defense perspective, with prior knowledge about the application domain (training set), it is also possible to finetune AT against specific semantic classes (vulnerable classes). Due to space, we leave this interesting problem to the future works.
@@
6 More examples of NMS and Latency Attacks
@@
7 When DETR meets Latency Attacks
@@
8 More Details of Experimental Settings
@@
9 Analysis of Training Process
@@ 10.1 Attack Strength
10.2 Balance between Precision/Recall
@@
11 Framework and Hardware Optimization under Latency Attacks
@@
In addition to PyTorch implementation, we also convert YOLOv5s and YOLOv8s to other implementations including ONNX and TensorRT for specialized acceleration. For reproducing purposes, the environment setup is shown in the Table 5. We employ the same method to attack the implementations of ONNX and TensorRT. In details, we export ONNX and TensorRT models using the official implementation, which convert only the backbone without utilizing the INMSLayer or EfficientNMSPlugin. Other configurations remain the same as 5.

### v2

# Can’t Slow me Down: Learning Robust and Hardware-Adaptive Object Detectors against Latency Attacks for Edge Devices

@@ Abstract
Object detection is a fundamental enabler for many real-time downstream applications such as autonomous driving, augmented reality and supply chain management. However, the algorithmic backbone of neural networks is brittle to imperceptible perturbations in the system inputs, which were generally known as misclassifying attacks. By targeting the real-time processing capability, a new class of latency attacks are reported recently. They exploit new attack surfaces in object detectors by creating a computational bottleneck in the post-processing module, that leads to cascading failure and puts the real-time downstream tasks at risks. In this work, we take an initial attempt to defend against this attack via background-attentive adversarial training that is also cognizant of the underlying hardware capabilities. We first draw system-level connections between latency attack and hardware capacity across heterogeneous GPU devices. Based on the particular adversarial behaviors, we utilize objectness loss as a proxy and build background attention into the adversarial training pipeline, and achieve a reasonable balance between clean and robust accuracy. The extensive experiments demonstrate the defense effectiveness of restoring real-time processing capability from $13$ FPS to $43$ FPS on Jetson Orin NX, with a better trade-off between the clean and robust accuracy. The source code is available at https://github.com/Hill-Wu-1998/underload
@@ 1 Introduction
Real-time object detection lies at the heart of a wide range of downstream applications such as autonomous driving [16], drone navigation [36] and video surveillance [20, 29]. From the early detectors such as (Faster) RCNN [9, 26] to the latest versions of the YOLO family [25, 33, 34], we have seen incredible achievements of performance-efficiency improvements, e.g., reaching 56-60% mAP on the MS-COCO benchmark with more than 30 FPS on embedded NVIDIA Jetson boards.
@@ 3.1 Algorithmic Vulnerability from NMS
$\small T_{\text{nms}}=\left\{\begin{aligned} T_{\text{base}},&~{}~{}~{}|C|\leq N_{t}\\ a|C|^{2},&~{}~{}~{}|C|>N_{t}\\ \end{aligned}\right.$ (1)
@@ 3.2 System-level Impacts of Latency Attacks
$\small T_{\text{nms}}=(\alpha\frac{|C|^{2}}{S_{\text{IoU}}}+\beta\frac{|C|}{B})<T-T_{\text{backbone}}\$ (3)
$\small\mathbb{E}[f_{\theta}(x)]_{\text{box}}<\frac{S_{\text{IoU}}}{2\alpha}\sqrt{\frac{\beta^{2}}{B^{2}}-4\frac{\alpha}{S_{\text{IoU}}}(T-T_{\text{backbone}})}$ (4)
@@ 4 Background-Attentive Adversarial Training
$\small\theta^{\ast}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{(x,y)\sim\mathbb{D}}\bigl{[}\max_{\left\lVert\delta\right\rVert_{p}\leq\epsilon}\mathcal{L}(f_{\theta}(x+\delta),y)\bigr{]}.$ (5)
@@ 4.1 Objectness Loss
$\small\theta$ $=$ $\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\mathbf{x}\sim\mathcal{D},\mathbf{y},\mathbf{b}}\mathcal{L}(f_{\theta}(\mathbf{x}),\mathbf{y},\mathbf{b})$ (6)
@@
$\vspace{-0.03in}\small\mathcal{L}_{obj}=\sum_{i=1}^{K}\big{[}\mathds{1}_{i}\ell_{BCE}(1,\hat{C}_{i})+(1-\mathds{1}_{i})\ell_{BCE}(0,\hat{C}_{i})\big{]}.\vspace{-0.03in}$ (7)
@@
An essential function of objectness is to differentiate objects from the background. Denote the objects as $x_{\text{obj}}=\bigcup_{k=1}^{K}(b_{i})^{\mathbf{y}}$, $x^{\prime}_{\text{obj}}=x_{\text{obj}}+\delta_{\text{obj}}$, and the background as $x_{\text{bg}}=x-x_{\text{obj}},x^{\prime}_{\text{bg}}=x_{\text{bg}}+\delta_{\text{bg}}$. We define background boundary margin as the distance for a background region to the decision boundary of becoming a “phantom object” in the pixel space $\mathcal{X}$ [39],
@@
$\small\mathbb{E}_{x\sim\mathbb{D}}[B_{\theta}(x^{\prime}_{\text{bg}})]\leq\mathbb{E}_{x\sim\mathbb{D}}[B_{\theta}(x^{\prime}_{\text{obj}})],$ (9)
@@ 4.2 Proposed AT Method
$\small\theta^{\ast}=\operatorname*{arg\,min}\limits_{\theta,{\delta^{M}\in\mathcal{S}}}\mathcal{L}\big{(}f_{\theta}(\mathbf{x}+\delta^{M}),\mathbf{b},\mathbf{y}\big{)}\vspace{-0.05in}$ (10)
@@
$\small\mathcal{S}=\bigl{\{}\delta^{M}\leftarrow~{}\mathbf{M}\odot\operatorname*{arg\,max}\limits_{\left\lVert\delta\right\rVert_{p}\leq\epsilon,i\in\mathcal{K}}\mathcal{L}_{obj}(f_{\theta}(\mathbf{x}+\delta),b_{i},y_{i})\bigr{\}}$ (11)
$\small\mathbb{E}[f_{\theta}(x+\delta^{M})]_{\text{box}}<\frac{S_{\text{IoU}}}{2\alpha}\sqrt{\frac{\beta^{2}}{B^{2}}-4\frac{\alpha}{S_{\text{IoU}}}(T-T_{\text{backbone}})}\$ (12)
$\small\mathbf{M}=\{\mathbf{0},\mathbf{1}\}^{r_{x}^{i}\times r_{y}^{i}},r_{x}^{i}\in[0,W_{b}^{i}],r_{y}^{i}\in[0,H_{b}^{i}],i\in\mathcal{K}.$ (13)
@@
Property 3 (Monotonicity). The box count $\mathbb{E}[f_{\theta}(x+\delta^{M})]_{\text{box}}$ is monotonously increasing regarding the mask size $r_{x}^{i}\times r_{y}^{i}$.
@@ 5.5 Background and Object Semantics
Finally, we expose intriguing artifacts between the background and object semantics through the lens of targeted attacks, that echo with our design of background-attentive AT. The goal is to launch targeted attacks to generate phantom objects of “cars” and “boats” under different background regions to different images in Fig. 8. For #1, it is much easier to generate more “cars” on the track than “boats” and the opposite is true for #2. For #3, it is a bit easier to generate more “boats” in the sky than “cars”, due to the blue color of the sky similar to the ocean. For #4, it is equally possible to generate as many “cars” or “boats” since the blank image contains no semantic information.
From the attacker’s perspective, the attack success rate can be improved by targeting the closest semantics to the background region, E.g., generating phantom “birds” in the “sky”, or “persons” on the “road”, which not only requires less perturbation (stealthier), but also generates more phantoms with higher latency in the NMS. This is because the semantically close objects/background are more vulnerable to adversarial attacks that slightly push the objects/background region over the manifold to be adversarial. From the defense perspective, with prior knowledge about the application domain (training set), it is also possible to finetune AT against specific semantic classes (vulnerable classes).
6 Conclusion
Acknowledgment. This work is supported in part by NSF of Zhejiang Province LZ25F020007, NSFC 62394341, 92467301, 62293511, the Fundamental Research Funds for the Central Universities 226202400182 and the ZJUCSE-Enflame Cloud and Edge Intelligence Joint Laboratory.
Supplementary Material
@@ 6 Conclusion
A More examples of NMS and Latency Attacks
@@
B When DETR meets Latency Attacks
@@
C More Details of Experimental Settings
@@
D Analysis of Training Process
@@ E.1 Attack Strength
E.2 Balance between Precision/Recall
@@
F Framework and Hardware Optimization under Latency Attacks
@@
In addition to PyTorch implementation, we also convert YOLOv5s and YOLOv8s to other implementations including ONNX and TensorRT for specialized acceleration. For reproducing purposes, the environment setup is shown in the Table 5. We employ the same method to attack the implementations of ONNX and TensorRT. In details, we export ONNX and TensorRT models using the official implementation, which convert only the backbone without utilizing the INMSLayer or EfficientNMSPlugin. Other configurations remain the same as Sec. 5.